OpenStudy (anonymous):

can someone graph this equation y=(1/10)^x

OpenStudy (anonymous):

What level of math are you are at, are you plotting points or using graphing calculator?

OpenStudy (anonymous):

i am in 7th grade and i ne how to do this lol

OpenStudy (anonymous):

will is

OpenStudy (radar):

I can get u started. when x=0, y=1, when x=1, y=.1, you better take it from there.

OpenStudy (anonymous):

y= (1^x) / (10^x) by distributing the power

OpenStudy (anonymous):

now , 1 to the power of anything is 1 so y= 1 / (10^x) bring the denominator up the top so y= 10^(-x)

OpenStudy (anonymous):

this is very similar to y=e^(-x) , so it will be large and positive in the top left hand section of the xy plane , then it will decay and approach the x axis as x-> positive infinity